how much energy is transferred by battery of end 4.5V and 1.0C of charge passes through it
1
Expert's answer
2019-09-30T10:12:08-0400

The potential difference is equal to the work per unit charge


That is to say: "V=\frac{E}{Q}"


Where

  • Potential difference. "V=4.5V"
  • Energy "E=??"
  • Electric charge "Q=1.0C"


Expression for energy


"E=Q*V"


Numerically evaluating


"E=1.0C*4.5V=4.5J"


The transferred energy is "\boxed{E=4.5J} \text[ j=joule]"